OpenStudy (anonymous):

What is the nth term for the sequence= 5.5, 7, 8.5, 10, 11.5, ..... ?

OpenStudy (saifoo.khan):

11.5 + 1.5 = answer.

OpenStudy (anonymous):

I am with saif !

OpenStudy (anonymous):

nonononn, not the next term, the Nth term.

OpenStudy (saifoo.khan):

do you mean the formula?

OpenStudy (anonymous):

yeahh.

OpenStudy (anonymous):

\[a_n = 5.5 + n(1.5)\]

OpenStudy (saifoo.khan):

Righto^

OpenStudy (anonymous):

no, how will it work for the first one? n is one there/

OpenStudy (anonymous):

\[a_n = 5.5 + (n-1)(1.5)\]

OpenStudy (anonymous):

ohhhhh, thanks :D